OpenStudy (anonymous):

Write an equation in point-slope form for a horizontal line that passes through (4, -2)

OpenStudy (anonymous):

The point-slope equation of the line is: y−y1=m(x−x1) Substituting the values x1=4 and y1=−2 and as it is a horizontal line m=0 y−(−2)=0(x−4) y+2=0 Finally, the ordinary equation of the line is: y=−2